OpenStudy (anonymous):

If line AB contains the points (6, -2) and (-3, 16), what is the slope of a line perpendicular to line AB ? Enter your answer as a decimal.

OpenStudy (anonymous):

Do you know how to find the slope of those 2 points?

OpenStudy (anonymous):

m= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

Yeah I do, just double checking.

OpenStudy (anonymous):

m=(16+2)/(-3-6) m=(18)/(-9) m=-2

OpenStudy (anonymous):

Okay thanks

OpenStudy (anonymous):

then to find the perpendicular slope, you have to take the negative reciprical of that so the perpendicular slope is 1/2

OpenStudy (anonymous):

as a decimal, the answer is 0.5